OpenStudy (princeevee):

what is the hieght of a cone that has a diameter of 12 cm. and a lenght of 10 cm? ((will medal)))

OpenStudy (anonymous):

what are the answers to this question?

OpenStudy (princeevee):

you have to guess it or somethin its not multiple choice : (

OpenStudy (anonymous):

oh so you have to type in the answer?

OpenStudy (princeevee):

yeah

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

Do you go to FLVS????

OpenStudy (anonymous):

do you do your work in connections academy?

OpenStudy (anonymous):

uh no i dont know what that is

OpenStudy (princeevee):

the link isnt showing for me

OpenStudy (anonymous):

Florida Virtual School.

OpenStudy (princeevee):

?

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

one second i cant add the attachment

OpenStudy (anonymous):

Find the volume of a cone the radius of whose base is 21 cm and height is 28 cm. Solution : r = 21 cm and h = 28 cm Volume of cone = 1/3 π r 2 h V = 1/3 ( 3.14 x 21 x 21 x 28) V = 1/3 x 38772.72 ∴ Volume of a cone = 12924.24 cm 3

OpenStudy (anonymous):

there

OpenStudy (princeevee):

is this the question?

OpenStudy (anonymous):

I kinda understand how to do this!

OpenStudy (princeevee):

im looking for the hieght of the cone